OpenStudy (anonymous):

using the elimination method to solve the system of equations 5u + 3v = -17 2u + v = -5 need to find the ordered pair

OpenStudy (anonymous):

multiply the second equation by 3 and subtract from the first equation.wat do u get?

OpenStudy (anonymous):

u = 1

OpenStudy (anonymous):

no check ur calculation u=2

OpenStudy (anonymous):

okay I forgot to add the 1 u (5-6) on the other side

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

then would i find v by substituting 2*(2)+v=-5

OpenStudy (anonymous):

thats right.

OpenStudy (anonymous):

the ordered pair would be (2,3)

OpenStudy (anonymous):

nop check again

OpenStudy (anonymous):

are trying to find the ans?

OpenStudy (anonymous):

yes, would I multiply 2*(2)=4+3=7??

OpenStudy (anonymous):

would y = 4

OpenStudy (anonymous):

its not like that 2*(2)+v=-5 4+v=-5 v=-5-4 v=-9 got it?

OpenStudy (anonymous):

yes. thanks

OpenStudy (anonymous):

:)

OpenStudy (fellowroot):